OpenStudy (anonymous):

How many molecules are in each sample? 3.6gH2O

OpenStudy (abb0t):

Start by converting your sample (grams) to moles. Use the relationship of molar mass, which means grams per mol to convert to moles (as I stated previusly). Once you have moles, use avogrado's number to convert from moles to molecules. Avogadro's number is: \[6.022 \times 10^{23}\frac{ molecules }{mole }\]

OpenStudy (chmvijay):

18g(1mole) of water consists of = 6.023*10^23 molecules then 3.6 gof water consists of =6.023*10^23 *3.6 / 18 =

OpenStudy (chmvijay):

hey are you getting it or no !!! LOl:)!!!

OpenStudy (chmvijay):

any problem in understanding ?

OpenStudy (abb0t):

I don't understand.

OpenStudy (chmvijay):

what you didnt understood lol u only answered first no!!! iam asking to @bobvilla84

OpenStudy (anonymous):

sorry thought I logged off lol, but I got the answer 1.2044 * 10^23, ifit was supposed to be put in two significant figures, would it be 1.2?

OpenStudy (chmvijay):

LOL:)

OpenStudy (anonymous):

thank you!